# Completeness of coalition logics with seriality, independence of agents, or determinism

### v1

@@ 3.3.2 The lemma
      Note $\mathtt{impeach}({\sigma_{\mathtt{AG}}})=\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathtt{AG}\}\bmod n=\Big{(}\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathrm{B}_{j}\}+{{\sigma_{\mathtt{AG}}}(\overline{b})}\Big{)}\bmod n$. It suffices to show (($\bigg{(}\Big{(}\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathrm{B}_{j}\}\bmod n\Big{)}+\Big{(}{{\sigma_{\mathtt{AG}}}(\overline{b})}\bmod n\Big{)}\bigg{)}\bmod n=j$.
      Assume $j\geq\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})$. ($\Big{(}\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathrm{B}_{j}\}\bmod n\Big{)}+\Big{(}(j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}}))$ $\bmod n\Big{)}=\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+(j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}}))=j$. $j\bmod n=j$. Thus, $\mathtt{impeach}({\sigma_{\mathtt{AG}}})$ $=j$.
      Assume $j<\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})$. ($\Big{(}\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathrm{B}_{j}\}\bmod n\Big{)}+\Big{(}(j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+n)\bmod n\Big{)}=\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+(j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+n)=j+n$. $j+n\bmod n=j$. Then, $\mathtt{impeach}({\sigma_{\mathtt{AG}}})=j$.
@@ 3.4 Upward derivability lemma
Let $\mathsf{X}\in\mathrm{ES}$ and $\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$ be a standard disjunction.
Then, $\vdash_{{\mathsf{MCL}+\mathsf{X}}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$, if the following condition, called the $\mathsf{X}$-derivability-reduction-condition, is met:
@@
Assume $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma$. It is easy to see $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
@@
    Assume $\mathsf{D}\notin\mathsf{X}$. Then, there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to\psi_{j}$. By Rule $\mathtt{Mon}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. We respectively consider whether $\mathsf{S}$ is in $\mathsf{X}$ and whether $\mathsf{I}$ is in $\mathsf{X}$.
@@
      Assume $\mathsf{S}\notin\mathsf{X}$ and $\mathsf{I}\notin\mathsf{X}$. Note in this case, $\mathbb{I}^{\prime}$ is $\epsilon$-neat: $\mathbb{I}^{\prime}$ is not empty, and for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $\mathrm{A}_{i}\neq\emptyset$ and $\mathrm{A}_{i^{\prime}}\neq\emptyset$, then $i=i^{\prime}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{SIA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
@@
      Assume $\mathsf{S}\notin\mathsf{X}$ and $\mathsf{I}\in\mathsf{X}$. Note in this case, Axiom $\mathtt{A}\text{-}\mathtt{IA}$ is an axiom of ${\mathsf{MCL}+\mathsf{X}}$, and $\mathbb{I}^{\prime}$ is $\mathsf{I}$-neat: $\mathbb{I}^{\prime}$ is not empty, and for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $i\neq i^{\prime}$, then $\mathrm{A}_{i}\cap\mathrm{A}_{i^{\prime}}=\emptyset$. By Axiom $\mathtt{A}\text{-}\mathtt{IA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
@@
      Assume $\mathbb{I}^{\prime}$ is empty. Note in this case, $\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$ is $\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, we have $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
      Assume $\mathbb{I}^{\prime}$ is not empty. By Axiom $\mathtt{A}\text{-}\mathtt{SIA}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$ $\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.

### v2

@@ 3.5 Downward validity lemma
    Note $\mathtt{impeach}({\sigma_{\mathtt{AG}}})=\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathtt{AG}\}\bmod n=\Big{(}\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathrm{B}_{j}\}+{{\sigma_{\mathtt{AG}}}(\overline{b})}\Big{)}\bmod n$. It suffices to show $\bigg{(}\Big{(}\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathrm{B}_{j}\}\bmod n\Big{)}+\Big{(}{{\sigma_{\mathtt{AG}}}(\overline{b})}\bmod n\Big{)}\bigg{)}\bmod n=j$.
    Assume $j\geq\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})$. $\Big{(}\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathrm{B}_{j}\}\bmod n\Big{)}+\Big{(}(j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}}))$ $\bmod n\Big{)}=\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+(j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}}))=j$. $j\bmod n=j$. Thus, $\mathtt{impeach}({\sigma_{\mathtt{AG}}})$ =j=j.
    Assume $j<\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})$. $\Big{(}\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathrm{B}_{j}\}\bmod n\Big{)}+\Big{(}(j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+n)\bmod n\Big{)}=\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+(j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+n)=j+n$. $j+n\bmod n=j$. Then, $\mathtt{impeach}({\sigma_{\mathtt{AG}}})=j$.
@@ 3.6 Upward derivability lemma
Let $\mathsf{X}\in\mathrm{ES}$ and $\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$ be a standard disjunction.
Then, $\vdash_{{\mathsf{MCL}+\mathsf{X}}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$, if the following condition, called the $\mathsf{X}$-derivability-reduction condition, is met:
@@
Assume $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma$. It is easy to see $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.
@@
    Assume $\mathsf{D}\notin\mathsf{X}$. Then, there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to\psi_{j}$. By Rule $\mathtt{Mon}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. We respectively consider whether $\mathsf{S}$ is in $\mathsf{X}$ and whether $\mathsf{I}$ is in $\mathsf{X}$.
@@
      Assume $\mathsf{S}\notin\mathsf{X}$ and $\mathsf{I}\notin\mathsf{X}$. Note in this case, $\mathbb{I}^{\prime}$ is $\epsilon$-neat: $\mathbb{I}^{\prime}$ is not empty, and for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $\mathrm{A}_{i}\neq\emptyset$ and $\mathrm{A}_{i^{\prime}}\neq\emptyset$, then $i=i^{\prime}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{SIA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.
@@
      Assume $\mathsf{S}\notin\mathsf{X}$ and $\mathsf{I}\in\mathsf{X}$. Note in this case, Axiom $\mathtt{A}\text{-}\mathtt{IA}$ is an axiom of $\mathsf{MCL}+\mathsf{X}$}, and $\mathbb{I}^{\prime}$ is $\mathsf{I}$-neat: $\mathbb{I}^{\prime}$ is not empty, and for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $i\neq i^{\prime}$, then $\mathrm{A}_{i}\cap\mathrm{A}_{i^{\prime}}=\emptyset$. By Axiom $\mathtt{A}\text{-}\mathtt{IA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.
@@
      Assume $\mathbb{I}^{\prime}$ is empty. Note in this case, $\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$ is $\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\emptyset]\hskip-2.0pt\rangle\top\rightarrow\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, we have $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.
      Assume $\mathbb{I}^{\prime}$ is not empty. By Axiom $\mathtt{A}\text{-}\mathtt{SIA}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\langle\hskip-2.0pt[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\hskip-2.0pt\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}$ $\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle\hskip-2.0pt[\mathrm{A}_{i}]\hskip-2.0pt\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle\hskip-2.0pt[\mathrm{B}_{j}]\hskip-2.0pt\rangle\psi_{j})$.